OpenStudy (anonymous):

Darian gets paid $6.35 an hour to help out at his dad's shop. He also mows his neighbor's lawn each week for $25. He earned $119.85 in the last two weeks. How many hours did Darian work at his dad's shop in those two weeks?

OpenStudy (anonymous):

Work backwards: 119.85 minus the two weeks(25*2) at the neighboors: 119.85 - 50 = 69.85 Now divide that by the amount paid an hour: 69.85 /6.35 = 11 so he worked for 11 hours at his dads shop:)

OpenStudy (ganpat):

6.35 x + 2 (25) = 119.85 6.35x + 50 = 119. 85 6.35x = 69.85 so, x = 11 hrs.. so he worked for 11 hrs..

OpenStudy (anonymous):

^^That's really how your supposed to figure it out in algebra so thumbs up to @Ganpat d(*_*)b

OpenStudy (ganpat):

@jazy : Thanks man :)

OpenStudy (anonymous):

im a girl lol np:)

OpenStudy (ganpat):

oops, my bad !!! :-p